Dunston Military Academy has an annual deficit of $250,000. Its 1,000 students pay tuition of $10,000 each per year. The economics faculty has recommended solving the problem by recruiting additional athletes with $5,000 scholarships. Each additional athlete will cost the school $2,500 (equipment, etc.). Assuming the academy agrees, how many athletes are needed to eliminate the deficit?

a.

25

b.

100

c.

50

d.

Zero, the deficit cannot be eliminated by giving more scholarships.
Answer #1

Answer

Cost of a student =2500

Fees =$10000 scholarship =5000

Gain from one student =10000-5000=500

Net gain =5000-2500=2500

number of students required =deficit /net gain per student

=250000/2500

=100

It will require 100 students

Option b